A contractor has to supply 10000 bearings per day to an automobile manufacturer. He finds that when he starts production run, he can produce 25000 bearings per day. The cost of holding a bearing in stock for one year is $2 and the set-up cost of a production run is $180. Find the economic order quantity(EOQ). How frequently should the production run be made?


1
Expert's answer
2021-07-30T13:52:17-0400

"\text{The formula for EOQ is given by $Q = \sqrt{\frac{2DS}{C}\frac{P}{P-d}}$}\\D=10000(365)=3650000\\d=10000\\S=180\\P=25000\\\implies\sqrt{\frac{2(3650000)(180)}{2}\frac{25000}{25000-10000}}=33090.78
\\\implies Q=33090.78\\\text{The frequency of production runs, f, is}\\f=\frac{Q}{d}=3.309"